PITMinutes 08-01-2016

From CASA Wiki
Jump to: navigation, search


R. Bachimanchi, E. Daly, K. Davis, C. Hovater, A. Hutton, G. Krafft, A. Kimber, C. Mounts, R. Nelson, T. Plawski, T. Powers, T. Reilly, C. Reece, H. Wang

Agenda and Notes

1. Operations Status and Summary

Enhanced Radiation monitoring in 1L26 (e-log benesch) systematic measurements started

Powers Strike Movie of Module Oscillation

Is there benefit from some kind of 40 Hz thumper measurment?

2. C50-12 Warm Girder Studies

Collected 60 samples from inside warm girders.

Secured and will be analyzed in two stages: gross and fine, regarding size and elements present

NL-12 cryomodule lift out tomorrow and will be delivered to SRF Inst.

3. C75 (Gigi/Bob Report Next Week)

Mock-up to be tested this week.

First look at HOM Qs and Q0 vs. E

C75 or C50 next year? Brief discussion. As not approved by Mont, looks like need to wait till 13th refurbished module. Need to still complete data gathering for future C75

4. Follow up Field emission Workshop (Geng)

Tentative Date September 12. Whole afternoon Subscribed. This room.

Preliminay agenda

Tentative Agenda

Action Items